the sum of two numbers is -3. if the second number is subtracted from the first, the difference is 27. find the numbers
Found 2 solutions by ankor@dixie-net.com, Mathtut:
Answer by ankor@dixie-net.com(12701)

the sum of two numbers is -3. if the second number is subtracted from the first, the difference is 27. find the numbers
:
x + y = -3
x - y = 27
----------------addition eliminates y, find x
2x = 24
x = 12
:
Find x using the 1st equation
12 + y = -3
y = -3 - 12
y = -15
:
Check solutions in the 2nd equation
12 - (-15) = 27

Answer by Mathtut(3670)

so lets call the 2 numbers a and b. a+b= -3 and a-b= -27 now solve by substitution or similtaneously. I choose substitution so solving for the first equation a=-3-b substituting this into the second equation we get -3-b-b=-27 adding 3 to both sides and combining like b terms gives us -2b=-24 dividing both sides by -2 gives us b=12 no plugging the value of b into either equation lets take the first a+12= -3 after subtracting 12 from both sides we get a= -15
so a= -15
b= 12
checking by substituting these back into the original equations we see that we have found the correct answers.